OpenStudy (anonymous):

the quantity of x plus 4 over 3 = 2

OpenStudy (anonymous):

\[{x+4\over3}=2\]solve

OpenStudy (anonymous):

(x+4)/3=2 i.e. x+4 = 6 i.e. x= 6-4 i.e. x= 2